OpenStudy (anonymous):

if f(x)=cos(x) and g(x)=tan(x) which of these lines is an asymptote of the graph of (f-g)(x)? a) x=π/4 b) x=π/2 c) x=π d) x=2π

OpenStudy (anonymous):

tangent is undefined at \(x=\frac{\pi}{2}\)

OpenStudy (anonymous):

soo..... would it be x=π/2 ??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

oh okay.. thanks!

OpenStudy (anonymous):

yw